OpenStudy (anonymous):

Gabriel measured the length of two icicles hanging from his porch. The first icicle was 29 centimeters 2 millimeters and the second icicle was 15 centimeters 6 millimeters. How much longer is the first icicle than the second icicle?

OpenStudy (bahrom7893):

convert it all to mm. How many mms are in a cm?

OpenStudy (bahrom7893):

1cm=10mm

OpenStudy (bahrom7893):

So 29cm=29*10mm=290mm

OpenStudy (bahrom7893):

So total length of the first icicle is 290+2=292 mm

OpenStudy (bahrom7893):

Repeat for the second one and you'll get 156

OpenStudy (anonymous):

its cm and mm not just mm

OpenStudy (bahrom7893):

i just converted everything to mm

OpenStudy (bahrom7893):

29cm and 2mm = 292 mm, please read my explanation, I'm not just going to post the answers. Anyway 292-156 = 136mm.

OpenStudy (bahrom7893):

You can convert that back to cm and mm

OpenStudy (anonymous):

13 centimeters 6 millimeters 13 centimeters 8 millimeters 14 centimeters 4 millimeters 14 centimeters 6 millimeters thats a b c and d in order

OpenStudy (bahrom7893):

like i said im not just gonna give u the answer.

OpenStudy (anonymous):

but i dont know how to do this

OpenStudy (anonymous):

someone help meee!!!!!!!

OpenStudy (anonymous):

can u answer this

OpenStudy (bahrom7893):

come on man i even gave u the answer.. 136mm

OpenStudy (bahrom7893):

Convert it back to cm and mm

OpenStudy (bahrom7893):

there are 10 mm in 1 cm

OpenStudy (anonymous):

ok here i go...

OpenStudy (bahrom7893):

ok.. post ur steps

OpenStudy (anonymous):

136 mm *10=1360

OpenStudy (anonymous):

making it 13 cm and 6 mm?

OpenStudy (anonymous):

i got it right thank you for your time nice sir

OpenStudy (bahrom7893):

well that's the answer but ur steps were wrong, you should have divided it by 20

OpenStudy (bahrom7893):

i meant divided it by 10